its summer and i forgot how to do these kind of problems please help me understand it, the area of a rectangle is 48 sq ft. find all side lengths. the length is 3x and width is 4x
Answer by rfer(12670)

3x*4x=48
12x^2=48
x^2=4
x=2
--------------
3*2*4*2=48